OpenStudy (anonymous):

rewrite with rational exponents. (sqrt10mn)^3.

jhonyy9 (jhonyy9):

\[10mn ^{3/2}\]

jimthompson5910 (jim_thompson5910):

\[\Large \left(\sqrt{10mn}\right)^3\] \[\Large \left((10mn)^{\frac{1}{2}}\right)^3\] \[\Large (10mn)^{\left(\frac{1}{2}\right)\times3}\] \[\Large (10mn)^{\frac{3}{2}}\] So \[\Large \left(\sqrt{10mn}\right)^3 = (10mn)^{\frac{3}{2}}\]